OpenStudy (anonymous):

Amy sold pens and pencils for a fundraiser at school. Pencils sold for $1.50 and pens sold for $2. She sold a total of 35 items and made $61.50. How many pens did Amy sell?

hero (hero):

x + y = 35 1.50x + 2y = 61.50

hero (hero):

y = 35 - x y = (61.50 - 1.50x)/2 y = y 2(35 - x) = 61.50 - 1.50x 70 - 2x = 61.50 - 1.50x 70 - 61.50 = 2x - 1.50x 8.50 = .50x 8.50/.50 = x 17 = x 18 = y